# Small-time asymptotics and the emergence of complex singularities for the KdV equation

### v1

@@ C.2 Solving for $G_{m}(\tau)$
We need to substitute the asymptotic expansion (68) into PII, $F^{\prime\prime}=2F^{3}+\xi F-\alpha$, to derive ode problems for the $G_{m}(\tau)$, $m\geq 0$. By applying the chain rule, we find
@@
$F^{\prime\prime}\sim$ $\,\,\xi^{3/2}\left(\tau^{2}G_{0}^{\prime\prime}+\tau G_{0}^{\prime}\right)+\tau^{2}G_{1}^{\prime\prime}+\tau G_{1}^{\prime}-{\textstyle\frac{3}{2}}\tau^{2}G_{0}^{\prime\prime}$
$\,+\sum_{m=0}^{\infty}\frac{1}{\xi^{3/2+3m/2}}\bigg{(}\tau^{2}G_{m+2}^{\prime\prime}+\tau G_{m+2}^{\prime}-{\textstyle\frac{3}{2}}\tau^{2}G_{m+1}^{\prime\prime}-3(1+m)\tau G_{m+1}^{\prime}$
$\hskip 34.44443pt+{\textstyle\frac{9}{16}}\tau^{2}G_{m}^{\prime\prime}+{\textstyle\frac{9}{16}}(1+4m)\tau G_{m}^{\prime}+\frac{9m^{2}-1}{4}G_{m}\bigg{)}$
$F^{3}\sim$ $\,\,\xi^{3/2}G_{0}^{3}+3G_{0}^{2}G_{1}+\sum_{m=0}^{\infty}\frac{1}{\xi^{3/2+3m/2}}\sum_{i=0}^{m+2}\sum_{j=0}^{m+2-i}G_{i}G_{j}G_{m+2-i-j},$ (128)
@@
$\tau^{2}G_{0}^{\prime\prime}+\tau G_{0}^{\prime}=2G_{0}^{3}+G_{0},$ (129)
$\tau^{2}G_{1}^{\prime\prime}+\tau G_{1}^{\prime}-{\textstyle\frac{3}{2}}\tau^{2}G_{0}^{\prime\prime}=6G_{0}^{2}G_{1}+G_{1}-\alpha,$ (130)
while analogous (but more complicated) odes for $G_{m}$, $m\geq 2$, can be also be derived using the information above. For boundary conditions, we apply (69) for $m=0$ and $1$, to give
$G_{0}\sim$   $\,\,0+\tau+0\,\tau^{2}+{\textstyle\frac{1}{4}}\tau^{3}+0\,\tau^{4}+\ldots$ (131)
$G_{1}\sim$ $\,\,\alpha-\frac{96\alpha^{2}-5}{48}\tau+2\alpha\tau^{2}-\frac{96\alpha^{2}-23}{64}\tau^{3}+\tau^{4}+\ldots.$ (132)
Again, boundary conditions for other $G_{m}$ can be taken from (69).
Leaving out the details, we can solve (129)–(132) using Maple (or equivalent) to give
$G_{0}=\frac{4\tau}{4-\tau^{2}},\quad G_{1}=\frac{3\tau^{5}+24\alpha\tau^{4}+(118-192\alpha^{2})\tau^{3}+576\alpha\tau^{2}+(40-768\alpha^{2})\tau+384\alpha}{24(\tau^{2}-4)^{2}}.$ (133)
Thus, we have a better approximation than (70) for $F$ in the neighbourhood of the negative-$\xi$ axis in the far field, namely $F\sim\xi^{1/2}G_{0}(\tau)+G_{1}(\tau)/\xi$ with (133) and $\tau$ given by (66). For the next subsection, we shall use the results
$G_{0}\sim\frac{-2}{\tau+2},\quad G_{1}\sim\frac{8\alpha(\alpha+1)-35/12}{(\tau+2)^{2}}\quad\mbox{as}\quad\tau\rightarrow-2,$ (134)
$G_{0}\sim\frac{-2}{\tau-2},\quad G_{1}\sim\frac{-8\alpha(\alpha-1)+35/12}{(\tau-2)^{2}}\quad\mbox{as}\quad\tau\rightarrow 2.$ (135)
@@ C.3 Singularities near anti-Stokes lines and higher-order corrections
$\tau_{0}\sim\tau_{0}^{(0)}+\frac{\tau_{0}^{(1)}}{\xi^{3/2}}+\frac{\tau_{0}^{(2)}}{\xi^{3}}+\ldots\quad\mbox{as}\quad|\xi|\rightarrow\infty.$ (136)
From (133), there are two options, namely
@@
provided $\tau_{0}^{(1)}=k_{1}/k_{0}$. Thus, comparing with (134)–(135), we find the two-term approximations for the singularities in terms of $\tau$ come from (136) with either
@@
With these correction terms we can provide a more accurate approximation for the location of the singularities than (72). The updated version of (71) becomes
@@
$+\frac{1}{(3\pi n)^{4/3}}\left(-\frac{1}{4}\left(\frac{3\pi}{4}+\frac{\mathrm{i}}{2}\ln(3\pi n)+\mathrm{i}\log\left(\frac{\tau_{0}^{(0)}}{\sigma_{1}}\right)+\frac{3\mathrm{i}}{2}\right)^{2}-\frac{9}{16}+\frac{\tau_{0}^{(1)}}{\tau_{0}^{(0)}}\right)\quad\mbox{as}\quad n\rightarrow\infty,$ (137)

### v2

@@ C.2 Solving for $G_{m}(\tau)$
We need to substitute the asymptotic expansion (3.49) into PII, $F^{\prime\prime}=2F^{3}+\xi F-\alpha$, to derive ode problems for the $G_{m}(\tau)$, $m\geq 0$. By carefully applying the chain rule, we find
@@
$\frac{\mathrm{d}^{2}F}{\mathrm{d}\xi^{2}}\sim$ $\,\,\xi^{3/2}\left(\tau^{2}\frac{\mathrm{d}^{2}G_{0}}{\mathrm{d}\tau^{2}}+\tau\frac{\mathrm{d}G_{0}}{\mathrm{d}\tau}\right)+\tau^{2}\frac{\mathrm{d}^{2}G_{1}}{\mathrm{d}\tau^{2}}+\tau\frac{\mathrm{d}G_{1}}{\mathrm{d}\tau}-{\textstyle\frac{3}{2}}\tau^{2}\frac{\mathrm{d}^{2}G_{0}}{\mathrm{d}\tau^{2}}$
$\,+\sum_{m=0}^{\infty}\frac{1}{\xi^{3/2+3m/2}}\bigg(\tau^{2}\frac{\mathrm{d}^{2}G_{m+2}}{\mathrm{d}\tau^{2}}+\tau\frac{\mathrm{d}G_{m+2}}{\mathrm{d}\tau}-{\textstyle\frac{3}{2}}\tau^{2}\frac{\mathrm{d}^{2}G_{m+1}}{\mathrm{d}\tau^{2}}-3(1+m)\tau\frac{\mathrm{d}G_{m+1}}{\mathrm{d}\tau}$
$\hskip 43.05542pt+{\textstyle\frac{9}{16}}\tau^{2}\frac{\mathrm{d}^{2}G_{m}}{\mathrm{d}\tau^{2}}+{\textstyle\frac{9}{16}}(1+4m)\tau\frac{\mathrm{d}G_{m}}{\mathrm{d}\tau}+\frac{9m^{2}-1}{4}G_{m}\bigg)$
$F^{3}\sim$ $\,\,\xi^{3/2}G_{0}^{3}+3G_{0}^{2}G_{1}+\sum_{m=0}^{\infty}\frac{1}{\xi^{3/2+3m/2}}\sum_{i=0}^{m+2}\sum_{j=0}^{m+2-i}G_{i}G_{j}G_{m+2-i-j},$ (C.1)
@@
$\tau^{2}\frac{\mathrm{d}^{2}G_{0}}{\mathrm{d}\tau^{2}}+\tau\frac{\mathrm{d}G_{0}}{\mathrm{d}\tau}=2G_{0}^{3}+G_{0},$ (C.2)
$\tau^{2}\frac{\mathrm{d}^{2}G_{1}}{\mathrm{d}\tau^{2}}+\tau\frac{\mathrm{d}G_{1}}{\mathrm{d}\tau}-{\textstyle\frac{3}{2}}\tau^{2}\frac{\mathrm{d}^{2}G_{0}}{\mathrm{d}\tau^{2}}=6G_{0}^{2}G_{1}+G_{1}-\alpha,$ (C.3)
while analogous (but more complicated) odes for $G_{m}$, $m\geq 2$, can be also be derived using the information above. For boundary conditions, we apply (3.50) for $m=0$ and $1$, to give
$G_{0}\sim$   $\,\,0+\tau+0\,\tau^{2}+{\textstyle\frac{1}{4}}\tau^{3}+0\,\tau^{4}+\ldots$ (C.4)
$G_{1}\sim$ $\,\,\alpha-\frac{96\alpha^{2}-5}{48}\tau+2\alpha\tau^{2}-\frac{96\alpha^{2}-23}{64}\tau^{3}+\tau^{4}+\ldots.$ (C.5)
Again, boundary conditions for other $G_{m}$ can be taken from (3.50).
Leaving out the details, we can solve (C.2)–(C.5) using maple (or equivalent) to give
$G_{0}=\frac{4\tau}{4-\tau^{2}},\quad G_{1}=\frac{3\tau^{5}+24\alpha\tau^{4}+(118-192\alpha^{2})\tau^{3}+576\alpha\tau^{2}+(40-768\alpha^{2})\tau+384\alpha}{24(\tau^{2}-4)^{2}}.$ (C.6)
Thus, we have a better approximation than (3.51) for $F$ in the neighbourhood of the negative $\xi$-axis in the far field, namely $F\sim\xi^{1/2}G_{0}(\tau)+G_{1}(\tau)/\xi$ with (C.6) and $\tau$ given by (3.47). For the next subsection, we shall use the results
$G_{0}\sim\frac{-2}{\tau+2},\quad G_{1}\sim\frac{8\alpha(\alpha+1)-35/12}{(\tau+2)^{2}}\quad\mbox{as}\quad\tau\rightarrow-2,$ (C.7)
$G_{0}\sim\frac{-2}{\tau-2},\quad G_{1}\sim\frac{-8\alpha(\alpha-1)+35/12}{(\tau-2)^{2}}\quad\mbox{as}\quad\tau\rightarrow 2.$ (C.8)
@@ C.3 Singularities near anti-Stokes lines and higher-order corrections
$\tau_{0}\sim\tau_{0}^{(0)}+\frac{\tau_{0}^{(1)}}{\xi^{3/2}}+\frac{\tau_{0}^{(2)}}{\xi^{3}}+\ldots\quad\mbox{as}\quad|\xi|\rightarrow\infty.$ (C.9)
From (C.6), there are two options, namely
@@
provided $\tau_{0}^{(1)}=k_{1}/k_{0}$. Thus, comparing with (C.7)–(C.8), we find the two-term approximations for the singularities in terms of $\tau$ come from (C.9) with either
@@
With these correction terms we can provide a more accurate approximation for the location of the singularities than (3.53). The updated version of (3.52) becomes
@@
$+\frac{1}{(3\pi n)^{4/3}}\left(-\frac{1}{4}\left(\frac{3\pi}{4}+\frac{\mathrm{i}}{2}\ln(3\pi n)+\mathrm{i}\log\left(\frac{\tau_{0}^{(0)}}{\sigma_{1}}\right)+\frac{3\mathrm{i}}{2}\right)^{2}-\frac{9}{16}+\frac{\tau_{0}^{(1)}}{\tau_{0}^{(0)}}\right)\quad\mbox{as}\quad n\rightarrow\infty,$ (C.10)